# Orthogonal roots, quantum Hafnians, and generalized Rothe diagrams

### v1

@@ 5.1. $T(E_{8},8)$
The assumptions that $\lvert Q\cap Q_{0}\rvert=2<4$ and $\lvert Q\cap Q_{1}\rvert=2<4$ imply that $\mathrm{Res}^{(0)}(Q)=\emptyset$ and $\mathrm{Res}^{(2)}(Q)=\emptyset$ by Lemma 5.2 (ii) and (iii), respectively, and (i) follows.
The fact that $H\in D_{6}(R)$ follows from Lemma 2.14 (i). Let $\Xi=\text{\rm Span}(H)\cap\Phi$ be the corresponding $D_{6}$ subsystem in $\Phi$. We have $\theta\notin\Xi$ by Lemma 5.5 (i), so $\theta\notin\text{\rm Res}(H)$ and $\mathrm{Res}^{(2)}(H)=\emptyset$, proving the first equality in (ii). The two remaining equalities follow from Proposition 4.1 (i) and Lemma 5.2 (ii) in the same way as the last two equalities in Lemma 5.3 (ii). ∎
@@
Let $\beta\in\Xi\cap U_{1}$. If we have $\beta=\varepsilon_{1}-\varepsilon_{i}$ for some $2\leq i\leq 6$ or $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $i>a$, then we have $s_{\gamma}(\beta)<0$ for $\gamma=\varepsilon_{1}+\varepsilon_{a}\in H$, so $\beta\notin\text{\rm Res}(H)$. On the other hand, if $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $1<i<a$ then we have $s_{\gamma}(\beta)>0$ for all $\gamma\in H$ and hence $\beta\in\text{\rm Res}(H)$. It follows that $\mathrm{Res}^{(1)}(H)=\{\varepsilon_{1}+\varepsilon_{i}:1<i<a\}$ and $\rho^{({1})}(H)=a-2$. Similarly, we have $\rho^{({1})}(rH)=a^{\prime}-2$, so that $\rho^{({1})}(rH)-\rho^{({1})}(H)=a^{\prime}-a$. By Remarks 2.13 and 3.5 (ii), the assumption that $\rho(rQ)>\rho(Q)$ implies that we have one of the following situations:
@@
We have $\rho^{({1})}(rH)-\rho^{({1})}(H)=a^{\prime}-a\geq\rho(rQ)-\rho(Q)$ in all these cases, which proves (i). Now suppose that $\alpha$ is a simple root of $\Phi$. This implies that $\alpha$ is also a simple root of $\Xi$, so the reflection $r$ induces a basic transposition $s_{i}:=(i,i+1)$ on the perfect matching $M:=\{\{1,a\},\{b,c\}\}$ corresponding to $Q$, with $\{i,i+1\}\subseteq\{1,a,b,c\}$ because $Q$ is the support of $\alpha$. If $\{i,i+1\}\in M$, then $rH=H$ and (ii) follows. Otherwise, $s_{i}$ must move either $1$ or $a$ to an adjacent integer, and in either case the numbers $(b,c,d,e)$ and $(s_{i}(b),s_{i}(c),s_{i}(d),s_{i}(e))$ appearing in the matchings corresponding to $Q_{0}$ and $rQ_{0}$ must be in the same relative order. It follows that $\rho(Q_{0})=\rho(rQ_{0})$, proving (ii). ∎
@@
$\text{\rm Res}(R)=\text{\rm Res}(Q)\ \dot{\cup}\ \text{\rm Res}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}(\text{\rm Res}(H^{\prime})\setminus\text{\rm Res}(Q))$
@@
$\mathrm{Res}^{(1)}(R)=\mathrm{Res}^{(1)}(Q)\ \dot{\cup}\ \mathrm{Res}^{(1)}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}(\mathrm{Res}^{(1)}(H^{\prime})\setminus\text{\rm Res}(Q),$
where we have $\mathrm{Res}^{(1)}(Q)=\text{\rm Res}(Q)$ by Lemma 5.6 (i) and $\text{\rm Res}(Q)=\mathrm{Res}^{(1)}(Q)\subseteq\mathrm{Res}^{(1)}(H^{\prime})$ for all $Q\subseteq H^{\prime}\in D_{6}(R)$ by Proposition 4.1 (i). Taking cardinalities of the sets in the last equation then yields the equation
@@
by Lemma 5.7 (i). For any $Q\subseteq H^{\prime}\subseteq D_{6}(R)\setminus\{H\}$, we have $\lvert H^{\prime}\cap U_{0}\rvert=\lvert rH^{\prime}\cap U_{0}\rvert<4$ and therefore that $\mathrm{Res}^{(0)}(H^{\prime})=\mathrm{Res}^{(0)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (ii). Furthermore, we have either $H^{\prime}=Q_{1}\cup Q$ or $\lvert H^{\prime}\cap Q_{1}\rvert<4$. In the former case, Lemma 5.3 (ii) and Proposition 4.1 (i) imply that $\theta\in\text{\rm Res}(Q_{1})\subseteq\text{\rm Res}(H^{\prime})$ and similarly $\theta\in\text{\rm Res}(rQ_{1})\subseteq\text{\rm Res}(rH^{\prime})$, so we have $\mathrm{Res}^{(2)}(H^{\prime})=\mathrm{Res}^{(2)}(rH^{\prime})=\{\theta\}$. In the latter case, we have $\mathrm{Res}^{(2)}(H^{\prime})=\mathrm{Res}^{(2)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (iii). It follows that $\rho^{({0})}(H^{\prime})=\rho^{({0})}(rH^{\prime})$ and $\rho^{({2})}(H^{\prime})=\rho^{({2})}(H^{\prime})$ in all cases, so we have
@@
for any $Q\subseteq H^{\prime}\in D_{6}(R)\setminus\{H\}$, where the inequality holds by an application of Corollary 4.4 to the $6$-root $H^{\prime}$ of the $D_{6}$ system $\text{\rm Span}(H^{\prime})\cap\Phi$. Equation (5.5) now implies that $\rho^{({1})}(rR)>\rho^{({1})}(R)$, as desired. ∎
@@
    Using Remark 4.5 (ii) and Lemma 5.3 (i), one can show that exactly 135 of the positive $8$-roots in $\Omega_{\Phi_{+}}$ contain $\theta$ and that the remaining 1890 positive $8$-roots are in a 3-to-1 correspondence with the elements of the set $\Omega_{U}$ from Theorem 5.9. The 33-to-11 ratio comes from the fact the each element $Q_{1}\in\Omega_{U}$ can be extended in three ways to a positive $8$-root $R=Q_{1}\ \dot{\cup}\ Q_{0}$, where $Q_{0}$ is one of the three features in the $D_{4}$-subsystem given by $\Psi=\{\beta\in\Phi:B(\beta,\gamma)=0\text{ for all }\gamma\in Q_{1}\}$. Combining this fact with Lemma 5.3 (ii), Remark 3.5 (ii), and Remark 4.5 (ii) shows that the Poincaré polynomial of $\Omega_{U}$ is $PS_{U}(q)=[5]_{q}[9]_{q}[14]_{q}$, which agrees with Table 1. In particular, the set $\Omega_{U}$ contains $5\cdot 9\cdot 14=630$ triples.
@@
    Recall from Remark 4.5 (iii) that the $8$-root
@@
    is the unique $W(E_{8})$-minimal element of $\Omega_{\Phi_{+}}$ and has no residues with respect to $\Phi_{+}$. It follows from Theorem 4.3 that $s_{8}(R)$ has a single residue, namely, $\theta_{8}$. Lemma 5.3 (ii) then implies that the unique $W(E_{7})$-minimal element of $\Omega_{U}$ is the quadruple $\{s_{8}(\alpha_{7}),s_{8}(\theta_{6}),s_{8}(\theta_{7}),s_{8}(\theta)\}$.
@@ 5.2. $T(E_{7},7)$
In this subsection, we maintain the notation of Theorem 5.9, identify the root system $\Phi(E_{7})$ with the subsystem $\Phi^{\prime}=\Phi_{8,0}$ of $\Phi$ in the usual way, and identify the set $U^{\prime}=U(E_{7},7)$ with the subset of $\Phi^{\prime}_{+}$ consisting of roots of positive odd $7$-height. Let $\theta^{\prime}=\theta_{7}$ be the highest root of $\Phi^{\prime}$. The $7$-height of $\theta^{\prime}$ is 1 by Equation (2.5), so it follows that $U^{\prime}=\Phi^{\prime}_{7,1}$. We will use Theorem 5.9 to prove that $T(E_{7},7)$ is a QP-triple in this subsection. In other words, we will prove that $\Omega_{U^{\prime}}$ is a quasiparabolic set for $W_{I^{\prime}}\cong W(E_{6})$ with level function $\rho_{U^{\prime}}$, where $I^{\prime}=S\setminus\{s_{8},s_{7}\}$.
@@
    We have $\alpha_{8}\not\in\text{\rm Res}_{U}(Q)$.
@@
    Every element of $\text{\rm Res}_{U}(Q)$ is orthogonal to $\eta$.
@@
Let $\gamma\in\text{\rm Res}_{U}(Q)$. Remark 3.2 (i) implies that $\gamma\neq\eta$, so we have $B(\eta,\gamma)\neq 2$. If it were the case that $B(\eta,\gamma)=-1$, it would follow that $\eta+\gamma=\theta$ by considering $8$-heights. This would imply that $\gamma=\alpha_{8}$, contradicting (i). If we had $B(\eta,\gamma)=+1$, it would follow that $s_{\eta}(\gamma)<0$, and $\gamma$ would not be a residue. We conclude that $B(\eta,\gamma)=0$, and (ii) follows. ∎
@@
    For every $T\in\Omega_{U^{\prime}}$, we have a bijection $\text{\rm Res}_{U^{\prime}}(T)\rightarrow\text{\rm Res}_{U}(\iota(T))$ sending each $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$ to $s_{8}(\gamma)$. In particular, we have $\rho_{U^{\prime}}(T)=\rho_{U}(\iota(T))$.
@@
Let $T=\{\beta_{1},\beta_{2},\beta_{3}\}\in\Omega_{U^{\prime}}$ and $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$. Since $B(\eta,s_{8}(\gamma))=B(\theta,\gamma)=0$, we have $s_{s_{8}(\gamma)}(\eta)=\eta>0$. For each $i\in\{1,2,3\}$, we have $\gamma\neq\beta_{i}$ by Remark 3.2 (i), and we have $s_{\beta_{i}}(\gamma)>0$ since $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$, so (iv) implies that $s_{s_{8}(\beta_{i})}(s_{8}(\gamma))>0$. It follows that $s_{8}(\gamma)\in\text{\rm Res}_{U}(\iota(T))$.
Conversely, for any $\delta\in\text{\rm Res}_{U}(\iota(T))$, we have $\delta\in\eta_{U}^{\perp}$ by Lemma 5.11 (ii), so we have $\delta=s_{8}(\gamma)$ for some $\gamma\in U^{\prime}$ by (i). The assumption that $\delta\in\text{\rm Res}_{U}(\iota(T))$ implies that $\delta\notin\iota(T)$ and, therefore, that $\gamma\notin T$. Since $\delta=s_{8}(\gamma)\in\text{\rm Res}_{U}(\iota(T))$, it follows from (iv) that $\gamma\in\text{\rm Res}_{U^{\prime}}(T)$, which proves the first assertion of (v). The second assertion follows from the first by taking cardinalities. ∎
@@
    Recall from Remark 4.5 (ii) that there are $135$ positive $7$-roots in type $E_{7}$, and that the Poincaré polynomial of the associated to these 7-roots is $PS_{\Phi^{\prime}_{+}}(q)=[3]_{q}[5]_{q}[9]_{q}$. The $7$-roots are in a 3–to–1 correspondence with the elements of $\Omega_{U}^{\prime}$ similar to the correspondence mentioned in Remark 5.10 (i). One can thus use the correspondence to deduce that $PS_{U^{\prime}}(q)=[5]_{q}[9]_{q}$ and $\lvert\Omega_{U^{\prime}}\rvert=45$.

### v2

@@ 5.1. $T(E_{8},8)$
The assumptions that $\lvert Q\cap Q_{0}\rvert=2<4$ and $\lvert Q\cap Q_{1}\rvert=2<4$ imply that $\mathscr{D}^{(0)}(Q)=\emptyset$ and $\mathscr{D}^{(2)}(Q)=\emptyset$ by Lemma 5.2 (ii) and (iii), respectively, and (i) follows.
The fact that $H\in D_{6}(R)$ follows from Lemma 2.15 (i). Let $\Xi=\text{\rm Span}(H)\cap\Phi$ be the corresponding $D_{6}$ subsystem in $\Phi$. We have $\theta\notin\Xi$ by Lemma 5.5 (i), so $\theta\notin\mathscr{D}(H)$ and $\mathscr{D}^{(2)}(H)=\emptyset$, proving the first equality in (ii). The two remaining equalities follow from Proposition 4.1 (i) and Lemma 5.2 (ii) in the same way as the last two equalities in Lemma 5.3 (ii). ∎
@@
Let $\beta\in\Xi\cap U_{1}$. If we have $\beta=\varepsilon_{1}-\varepsilon_{i}$ for some $2\leq i\leq 6$ or $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $i>a$, then we have $s_{\gamma}(\beta)<0$ for $\gamma=\varepsilon_{1}+\varepsilon_{a}\in H$, so $\beta\notin\mathscr{D}(H)$. On the other hand, if $\beta=\varepsilon_{1}+\varepsilon_{i}$ for some $1<i<a$ then we have $s_{\gamma}(\beta)>0$ for all $\gamma\in H$ and hence $\beta\in\mathscr{D}(H)$. It follows that $\mathscr{D}^{(1)}(H)=\{\varepsilon_{1}+\varepsilon_{i}:1<i<a\}$ and $\rho^{({1})}(H)=a-2$. Similarly, we have $\rho^{({1})}(rH)=a^{\prime}-2$, so that $\rho^{({1})}(rH)-\rho^{({1})}(H)=a^{\prime}-a$. By Remarks 2.14 and 3.6 (ii), the assumption that $\rho(rQ)>\rho(Q)$ implies that we have one of the following situations:
@@
We have $\rho^{({1})}(rH)-\rho^{({1})}(H)=a^{\prime}-a\geq\rho(rQ)-\rho(Q)$ in all these cases, which proves (i). Now suppose that $\alpha$ is a simple root of $\Phi$. This implies that $\alpha$ is also a simple root of $\Xi$, so the reflection $r$ induces an adjacent transposition $s_{i}:=(i,i+1)$ on the perfect matching $M:=\{\{1,a\},\{b,c\}\}$ corresponding to $Q$, with $\{i,i+1\}\subseteq\{1,a,b,c\}$ because $Q$ is the support of $\alpha$. If $\{i,i+1\}\in M$, then $rH=H$ and (ii) follows. Otherwise, $s_{i}$ must move either $1$ or $a$ to an adjacent integer, and in either case the numbers $(b,c,d,e)$ and $(s_{i}(b),s_{i}(c),s_{i}(d),s_{i}(e))$ appearing in the matchings corresponding to $Q_{0}$ and $rQ_{0}$ must be in the same relative order. It follows that $\rho(Q_{0})=\rho(rQ_{0})$, proving (ii). ∎
@@
$\mathscr{D}(R)=\mathscr{D}(Q)\ \dot{\cup}\ \mathscr{D}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}(\mathscr{D}(H^{\prime})\setminus\mathscr{D}(Q))$
@@
$\mathscr{D}^{(1)}(R)=\mathscr{D}^{(1)}(Q)\ \dot{\cup}\ \mathscr{D}^{(1)}(R\setminus Q)\ \dot{\cup}\ \dot{\bigcup}_{Q\subseteq H^{\prime}\in D_{6}(R)}\mathscr{D}^{(1)}(H^{\prime})\setminus\mathscr{D}(Q),$
where we have $\mathscr{D}^{(1)}(Q)=\mathscr{D}(Q)$ by Lemma 5.6 (i) and $\mathscr{D}(Q)=\mathscr{D}^{(1)}(Q)\subseteq\mathscr{D}^{(1)}(H^{\prime})$ for all $Q\subseteq H^{\prime}\in D_{6}(R)$ by Proposition 4.1 (i). Taking cardinalities of the sets in the last equation then yields the equation
@@
by Lemma 5.7 (i). For any $Q\subseteq H^{\prime}\subseteq D_{6}(R)\setminus\{H\}$, we have $\lvert H^{\prime}\cap U_{0}\rvert=\lvert rH^{\prime}\cap U_{0}\rvert<4$ and therefore that $\mathscr{D}^{(0)}(H^{\prime})=\mathscr{D}^{(0)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (ii). Furthermore, we have either $H^{\prime}=Q_{1}\cup Q$ or $\lvert H^{\prime}\cap Q_{1}\rvert<4$. In the former case, Lemma 5.3 (ii) and Proposition 4.1 (i) imply that $\theta\in\mathscr{D}(Q_{1})\subseteq\mathscr{D}(H^{\prime})$ and similarly $\theta\in\mathscr{D}(rQ_{1})\subseteq\mathscr{D}(rH^{\prime})$, so we have $\mathscr{D}^{(2)}(H^{\prime})=\mathscr{D}^{(2)}(rH^{\prime})=\{\theta\}$. In the latter case, we have $\mathscr{D}^{(2)}(H^{\prime})=\mathscr{D}^{(2)}(rH^{\prime})=\emptyset$ by Lemma 5.2 (iii). It follows that $\rho^{({0})}(H^{\prime})=\rho^{({0})}(rH^{\prime})$ and $\rho^{({2})}(H^{\prime})=\rho^{({2})}(H^{\prime})$ in all cases, so we have
@@
for any $Q\subseteq H^{\prime}\in D_{6}(R)\setminus\{H\}$, where the inequality holds by an application of Corollary 4.5 to the $6$-root $H^{\prime}$ of the $D_{6}$-system $\text{\rm Span}(H^{\prime})\cap\Phi$. Equation (5.5) now implies that $\rho^{({1})}(rR)>\rho^{({1})}(R)$, as desired. ∎
@@
    By Example 4.13, the set $\Omega_{\Phi_{+}}$ contains exactly $2025$ positive $8$-roots, and exactly 135 of them contain $\theta$. One can show using Lemma 5.3 that the remaining 1890 positive $8$-roots are in a 3-to-1 correspondence with the elements of the set $\Omega_{U}$ from Theorem 5.9. The 33-to-11 ratio comes from the fact the each element $Q_{1}\in\Omega_{U}$ can be extended in three ways to a positive $8$-root $R=Q_{1}\ \dot{\cup}\ Q_{0}$, where $Q_{0}$ is one of the three features in the $D_{4}$-subsystem given by $\Psi=\{\beta\in\Phi:B(\beta,\gamma)=0\text{ for all }\gamma\in Q_{1}\}$. Combining this fact with Lemma 5.3 (ii), Remark 3.6 (ii), and Table 1 shows that the Poincaré polynomial of $\Omega_{U}$ is $PS_{U}(q)=[5]_{q}[9]_{q}[14]_{q}$, which agrees with Table 2. In particular, the set $\Omega_{U}$ contains $5\cdot 9\cdot 14=630$ triples.
@@
    Recall from Remark 4.4 (iv) that the $8$-root
@@
    is the unique $W(E_{8})$-minimal element of $\Omega_{\Phi_{+}}$ and has level $\rho(R)=0$. It follows from Theorem 4.3 that a single root dominates $s_{8}(R)$, namely, $\theta_{8}$. Lemma 5.3 (ii) then implies that the unique $W(E_{7})$-minimal element of $\Omega_{U}$ is the quadruple $\{s_{8}(\alpha_{7}),s_{8}(\theta_{6}),s_{8}(\theta_{7}),s_{8}(\theta)\}$.
@@ 5.2. $T(E_{7},7)$
In this subsection, we maintain the notation of Theorem 5.9, we identify the root system $\Phi(E_{7})$ with the subsystem $\Phi^{\prime}=\Phi_{8,0}$ of $\Phi$ in the usual way, and we identify the set $U^{\prime}=U(E_{7},7)$ with the subset of $\Phi^{\prime}_{+}$ consisting of roots of positive odd $7$-height. Let $\theta^{\prime}=\theta_{7}$ be the highest root of $\Phi^{\prime}$. The $7$-height of $\theta^{\prime}$ is 1 by Equation (2.4), so it follows that $U^{\prime}=\Phi^{\prime}_{7,1}$. We will use Theorem 5.9 to prove that $T(E_{7},7)$ is a QP-triple in this subsection. In other words, we will prove that $\Omega_{U^{\prime}}$ is a quasiparabolic set for $W_{I^{\prime}}\cong W(E_{6})$ with level function $\rho_{U^{\prime}}$, where $I^{\prime}=S\setminus\{s_{8},s_{7}\}$.
@@
    We have $\alpha_{8}\not\in\mathscr{D}_{U}(Q)$.
@@
    Every element of $\mathscr{D}_{U}(Q)$ is orthogonal to $\eta$.
@@
Let $\gamma\in\mathscr{D}_{U}(Q)$. Remark 3.3 (i) implies that $\gamma\neq\eta$, so we have $B(\eta,\gamma)\neq 2$. If it were the case that $B(\eta,\gamma)=-1$, it would follow that $\eta+\gamma=\theta$ by considering $8$-heights. This would imply that $\gamma=\alpha_{8}$, contradicting (i). If we had $B(\eta,\gamma)=+1$, it would follow that $s_{\eta}(\gamma)<0$, and $\gamma$ would not dominate $Q$. We conclude that $B(\eta,\gamma)=0$, and (ii) follows. ∎
@@
    For every $T\in\Omega_{U^{\prime}}$, we have a bijection $\mathscr{D}_{U^{\prime}}(T)\rightarrow\mathscr{D}_{U}(\iota(T))$ sending each $\gamma\in\mathscr{D}_{U^{\prime}}(T)$ to $s_{8}(\gamma)$. In particular, we have $\rho_{U^{\prime}}(T)=\rho_{U}(\iota(T))$.
@@
Let $T=\{\beta_{1},\beta_{2},\beta_{3}\}\in\Omega_{U^{\prime}}$ and $\gamma\in\mathscr{D}_{U^{\prime}}(T)$. Since $B(\eta,s_{8}(\gamma))=B(\theta,\gamma)=0$, we have $s_{s_{8}(\gamma)}(\eta)=\eta>0$. For each $i\in\{1,2,3\}$, we have $\gamma\neq\beta_{i}$ by Remark 3.3 (i), and we have $s_{\beta_{i}}(\gamma)>0$ since $\gamma\in\mathscr{D}_{U^{\prime}}(T)$, so (iv) implies that $s_{s_{8}(\beta_{i})}(s_{8}(\gamma))>0$. It follows that $s_{8}(\gamma)\in\mathscr{D}_{U}(\iota(T))$.
Conversely, for any $\delta\in\mathscr{D}_{U}(\iota(T))$, we have $\delta\in\eta_{U}^{\perp}$ by Lemma 5.11 (ii), so we have $\delta=s_{8}(\gamma)$ for some $\gamma\in U^{\prime}$ by (i). The assumption that $\delta\in\mathscr{D}_{U}(\iota(T))$ implies that $\delta\notin\iota(T)$ and, therefore, that $\gamma\notin T$. Since $\delta=s_{8}(\gamma)\in\mathscr{D}_{U}(\iota(T))$, it follows from (iv) that $\gamma\in\mathscr{D}_{U^{\prime}}(T)$, which proves the first assertion of (v). The second assertion follows from the first by taking cardinalities. ∎
@@
    Recall that by Remark 4.4 (ii), there are $135$ positive $7$-roots in type $E_{7}$, and the Poincaré polynomial of the associated to these 7-roots is $PS_{\Phi^{\prime}_{+}}(q)=[3]_{q}[5]_{q}[9]_{q}$. The $7$-roots are in a 3–to–1 correspondence with the elements of $\Omega_{U}^{\prime}$ similar to the correspondence mentioned in Remark 5.10 (i). One can thus use the correspondence to deduce that $PS_{U^{\prime}}(q)=[5]_{q}[9]_{q}$ and $\lvert\Omega_{U^{\prime}}\rvert=45$.